### v1

# Junsheng Huang1,2, Zhitao He1, Sandeep Polisetty3, Qingyun Wang3, May Fung1,2

Abstract
With the widespread application of large language models (LLMs), the issue of generating non-existing facts, known as hallucination, has garnered increasing attention. Previous research in enhancing LLM confidence estimation mainly focuses on the single problem setting. However, LLM awareness of its internal parameterized knowledge boundary under the more challenging multi-problem setting, which requires answering multiple problems accurately simultaneously, remains underexplored. To bridge this gap, we introduce a novel method, Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning), that separates the learning of answer prediction and confidence estimation during fine-tuning on instruction data. Extensive eperiments demonstrate that our method proposed outperforms baselines by up to 25% in average precision. [^1] Our code and resource will be released upon publication.
[Uncaptioned image]  MAC-Tuning: LLM Multi-Compositional Problem Reasoning with Enhanced Knowledge Boundary Awareness
Junsheng Huang1,2, Zhitao He1, Sandeep Polisetty3, Qingyun Wang3, May Fung1,2 1Hong Kong University of Science and Technology 2University of Illinois     3UMass Amherst jh103@illinois.edu      yrfung@ust.hk
1 Introduction
Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023), information retrieval (Ren et al., 2023), and recommendation systems (Liu et al., 2023). Yet, they often produce non-existing facts when faced with questions outside their parametric knowledge, which undermines their reliability (Maynez et al., 2020). Many efforts have been dedicated to mitigating LLM hallucination, such as leveraging knowledge boundaries to constrain the reasoning scope of LLMs to help them better distinguish between reliable and unreliable information Chen et al. (2024); Liang et al. (2024a); Zhang et al. (2024). Notably, these work mainly focus on the single-problem setting, where users repeatedly input questions and context for models to answer one by one.
Figure 1: An illustration of the multi-problem setting. Red indicates that the LLM’s output is inaccurate.
LLM hallucination in the multi-problem setting — in which a single input contains multiple distinct sub-questions for the model to extract and address — remains relatively underexplored. As seen in Figure 1, this is a fundamentally challenging setting because the language model must distinguish each sub-question, reason over different knowledge, and synthesize results cohesively. Undesirable overshadowing of context from one sub-question with another, and propagation of reasoning confusion, may compromise the reliability of LLMs in multi-problem answering (Cheng et al., 2023, Wang et al., 2024, Son et al., 2024). As LLM-based multi-problem reasoning becomes increasingly widespread due to its efficiency benefits in scenarios involving extensive shared contexts (e.g., task instructions, exemplars), reduced model access, and lower API costs, enhancing model confidence estimation calibration for this emerging class of reasoning demands growing attention and effort as well.
In this paper, we investigate the hallucinations in LLMs within the multi-problem setting and propose leveraging the knowledge boundary to simultaneously handle the composition of multiple problems. Inspired by Zhang et al. (2024), which advocates for encouraging the LLM to express confidence to reduce hallucinations, we introduce Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning) under multi-problem setting. Our approach involves several key steps. First, we identify the knowledge boundary between parametric knowledge and the multi-problem dataset to extract uncertain questions. Next, we automatically label the model’s confidence for both certain and uncertain data. These labeled data are then used to create multiple question-answer data and multiple QA-Confidence data so we can train the original model by separating the learning process of ground-truth answers and confidence, which enhances performance and reliability.
Our contributions can be summarized as follows:
  • We are the first to explore LLM confidence estimation under the more challenging multi-problem setting, where LLMs must handle multiple problems simultaneously.
  • We propose MAC-Tuning, which separates the learning process of answer and confidence predictions for enhancing knowledge boundary awareness and reducing hallucination.
  • Through extensive experiments across various settings and datasets, MAC-Tuning achieves an AP score gain of up to 25% over baselines in LLM multi-problem reasoning. Finally, we share our insights discovered to motivate future work.
2 Methodology
Figure 2 shows the data construction process for Multiple Answers and Confidence Stepwise Tuning (MAC-Tuning).
2.1 Multi-Problem Tuning Data Construction
First, we randomly combine n single problems from original datasets to construct our initial Multi-Problem dataset. We utilize this to compare LLMs’ outputs with ground-truth answers, for distinguishing the knowledge boundary between LLM parameters and instruction data. Specifically, for each individual problem in the multi-problem pair, we assign: “I am sure” if the model output aligns with ground-truth answer; “I am unsure” elsewise (e.g., Step 2 in Figure 2). With the assigned confidence labels, we construct Multi-Problem Tuning data as follows:
Multiple QA pair $D_{MultQA}$: We directly combine the questions and answers together, with Question $q_{i}$ as input and Answer $a_{i}$ as output label, to form $D_{MultQA}=[(q_{1},a_{1})...(q_{i},a_{i})...(q_{n},a_{n})]$.
Figure 2: We first construct the Multi-Problem dataset, and then use it to generate Multi-Problem Tuning data.
Multiple QA-Confidence pair $D_{MultQA,C}$: The input consists of an instruction for the LLM to express its confidence (i.e, certainty in correctness) for a given question-answer pair, while the output is the confidence level in linguistic form [^2] The template is in Appendix A.3.
2.2 Training and Inference
Using the Multi-Problem Tuning data, we conduct a two-step supervised fine-tuning process to train the model to answer questions and express confidence in a multi-problem setting. The objective for the first step, in answering question, is:
$\max_{\Theta_{0}}\sum_{(Q,A)\in D_{MultQA}}\log P(A|Q;\Theta_{0})$ (1)
The objective for the second step, in expressing confidence, is:
$\max_{\Theta_{1}}\sum_{(Q,A,C)\in D_{MultQA,C}}\log P(C|Q,A;\Theta_{1})$ (2)
where $Q$, $A$, and $C$ represent the sets of multiple questions, multiple answers, and multiple confidence levels, respectively. $\Theta_{0}$ and $\Theta_{1}$ represent the parameters of the base model and the model after the first step of fine-tuning, respectively.
Model Independent Sequential
CoQA ParaRel GSM MMLU MTI-Bench SQA
AP ECE AP ECE AP ECE AP ECE AP ECE AP ECE
LLaMA3 54.6 22.6 45.1 40.8 79.3 52.8 50.3 43.8 37.4 17.7 44.9 35.4
QA-Only 66.3 15.1 53.7 12.6 75.3 36.1 58.5 17.9 45.0 16.9 56.6 21.0
Single-QA 65.5 28.9 73.5 10.7 56.6 44.5 58.3 25.7 N/A N/A N/A N/A
Merge-AC 67.4 17.0 73.0 65.3 75.1 44.8 58.5 18.3 38.3 33.7 49.2 31.7
MAC-Tuning 69.8 7.33 76.1 3.61 79.9 3.16 63.1 12.5 64.0 13.4 65.0 14.6
Table 1: This is the confidence calibration result (%). We use one-shot CoT for GSM results. Bold font highlights the best performance for the dataset across different methods. We don’t apply Single-QA to the Sequential setting dataset, as doing so would disrupt the logical connections among the questions.
3 Experiment
3.1 Dataset
We validate the effectiveness of our method across different problem settings and datasets: for the Independent setting, where the questions are not related to each other, we use the CoQA (Reddy et al., 2019), GSM (Cobbe et al., 2021), MMLU (Hendrycks et al., 2021), and ParaRel (Elazar et al., 2021) datasets; for the Sequential setting, where the questions are logically related to each other, we use the MTI-Bench (Son et al., 2024) and SQA (Iyyer et al., 2017) datasets. These datasets are either Question Answer (QA) or Multiple Choice (MC) formats. Table 2 shows the details of the dataset. Further information on the distribution of certain and uncertain data among the training set across different datasets is detailed in Appendix A.4.
Independent Sequential
CoQA ParaRel GSM MMLU MTI-Bench SQA
Train 5006 7500 7468 2448 2400 3985
Test 5011 5584 1319 2439 600 925
Type QA QA QA MC QA QA
Table 2: Statistics of the datasets.
3.2 Evaluation Metrics
We directly compare the LLM generation to the ground-truth answer for the Question-Answer format. For Multiple-Choice format, we check the choice (A, B, C, D) and the option in the LLM generation. Across both types of answer generation tasks, we consider three evaluation metrics: (1) Average Precision (AP): We use AP to measure the precision in identifying and ranking relevant predictions. A higher AP score means the model has high certainty about correct answers and high uncertainty about wrong answers. (2) Expected Calibrated Error (ECE): We use ECE to measure how closely the predicted certainty reflects the true certainty of LLM Chen et al. (2023). Low ECE indicates better-calibrated predictions. (3) Accuracy: We compute accuracy as the fraction of correct responses amongst questions in which LLMs expressed certainty towards their answers.
3.3 Baselines
We compare MAC-Tuning with the base model and its variants in the multi-problem settings. We use LLaMA3-8B-Instruct (LLaMA3) (Dubey et al., 2024) as the backbone. For baseline QA-Only, we fine-tune the base model directly using the Multiple Question-Answer pairs to evaluate the effectiveness of the traditional instruction tuning method under the multi-problem setting. For baseline Single-QA, we use single-problem data to fine-tune and directly apply it to the multi-problem setting. For baseline Merge-AC, instead of separating the learning process of ground-truth answers and confidence, we directly let the model learn multiple answers along with their corresponding confidence levels [^3] Baseline examples are in Appendix A.8. Implementation details are in Appendix A.6..
3.4 Overall Performance
In Table 1, we report the results on multi-problem setting from three single questions combined together. MAC-Tuning achieves the best AP score across all datasets, showing up to a 15% improvement, along with a lower ECE. This suggests that after MAC-Tuning, the model becomes more adept at distinguishing between certain and uncertain questions, delivering more reliable results through improved confidence estimation in answer prediction. We also evaluate each model’s accuracy on every dataset. MAC-Tuning consistently outperforms the base model in accuracy by up to 45.8% and, on average, 23.7%. The reason is that we separate the tasks of learning correct answers and confidence within a multi-problem setting. After learning the ground-truth answer, the LLM can better understand confidence, while still retaining its ability to extract information, respond accurately, and address multiple problems simultaneously.
Ablation on Different Component
We further test three variants of the MAC-Tuning method in the multi-problem setting: QA-Only, which is MAC-Tuning without the confidence component; Single-QA, where we evaluate MAC-Tuning with single problem data; and Merge-AC, where we evaluate MAC-Tuning without separating the learning process of ground-truth answers and confidence. As seen from the results in Table 1, MAC-Tuning has up to 25% and, on average, 11% AP improvement compared with Merge-AC, reflecting that separating the learning process of ground-truth answers and confidence is crucial in multi-problem setting, as LLM cannot learn both in one time. The performance of Single-QA is better than the base model but worse than QA-Only in most cases, showing that LLM can aware the knowledge boundary under single-problem setting and transfer it to multi-problem setting, but it is not sufficient for LLM to answer multiple problems simultaneously.
3.5 Investigation on Out of Domain Settings
We perform MAC-Tuning on base model with Sequential setting dataset SQA and test it on other datasets, with the results as presented in Table 3. Even on out-of-domain datasets, MAC-Tuning still outperforms the base model, showing that it can effectively learn the multi-problem setting and generalize across different domains.
Metric CoQA Pararel MMLU MTI-Bench
Accuracy 59.3 70.3 52.6 57.8
AP score 62.2 58.7 53.8 81.7
ECE 10.4 9.64 8.95 16.1
Table 3: The result (%) for MAC-Tuning on SQA dataset and test on other datasets.
3.6 Analysis on Various Number of Questions
We explore different numbers of questions in the multi-problem setting to investigate how this varies the accuracy. We only do this for three Independent setting datasets, and the results are reported in Figure 3. MAC-Tuning consistently outperforms the base model in accuracy by at least 10.0% and, on average, 26.8%. For easy tasks like ParaRel, The ability of the base model to handle multiple problems simultaneously is even higher when compared with the traditional single-problem setting, indicating that LLM could leverage in-context learning and focus on relevant knowledge better under multi-problem setting. However, for other datasets like MMLU, MAC-Tuning performs slightly worse as the question number increases. A reasonable explanation is that it is out of the base model’s ability to learn too many hard tasks together but within effective scope to learn several easy tasks at the same time. Further studies are detailed in A.9.
Figure 3: Accuracy for combining different number (n) of single problem together. Solid lines represent MAC Tuning, while dashed lines represent LLaMA3.
3.7 Analysis on Different Base Model Choice
Table 4 shows the result from changing the base model to Qwen2-7B-Instruct (Yang et al., 2024). We observe that the performance trends remain consistent even with a different base model. MAC-Tuning continues to demonstrate an average precision (AP) gain of up to near 24% with a lower ECE, showcasing the effectiveness of learning ground-truth answers and confidence separately.
 Independent  Sequential
Approach   ParaRel    MMLU MTI-Bench        SQA
AP ECE AP ECE AP ECE AP ECE
Vanilla 54.3 37.8 68.1 25.3 48.8 31.3 30.3 54.6
MAC-Tuning 78.7 9.59 73.0 17.1 53.3 18.6 47.7 29.2
Table 4: Confidence calibration result (%) for Qwen2-7B-Instruct, with bold denoting the top performance across different methods.
4 Conclusion
In this paper, we introduce a novel method, MAC-Tuning, to enhance large language model (LLM) confidence calibration and reasoning robustness in the challenging yet underexplored multi-problem scenario. Our proposed approach automatically constructs multi-problem setting question-answer pairs with confidence annotations for identifying the intrinsic knowledge gap between parametric knowledge and instructional data. With this data constructed, we guide the LLM to better reason on answer prediction and confidence estimation separately, in multi-problem setting. Extensive experiments across different datasets show that our method significantly improves performance in areas where the original LLM struggles.
Limitation
While our work provides valuable insight on the new Multiple Question setting and introduces an innovative fine-tuning method, there are several limitations to acknowledge. First, although we experimented with various prompts, as is typical in prompt-based LLM studies, we cannot ensure that slight changes in prompts would not significantly alter the results. Second, due to constraints of cost, time, and computational resources, we selected a subset of experiments that we believe to be informative and representative. However, additional experiments across a wider range of datasets and LLMs might provide further insights. Lastly, in this new setting, there may be other underlying reasons for the experimental results. Future work will aim to address these limitations by expanding datasets and conducting new experiments to explore other potential factors affecting performance.
Appendix A Appendix
A.1 Full Case for Examples of Introduction
Full case for the examples in introduction can be found in Figure 4.
Figure 4: The full case of examples in introduction in Multiple Problem setting. Red context indicates that LLM’s output is inaccurate. The second answer lacks the information of "Pre-Vatican" and the third answer contains a completely factual error. After MAC-Tuning, LLM show uncertainty towards answering this two previously incorrect questions.
A.2 Related Work
Hallucination: Large language models (LLMs) are widely used in knowledge-intensive scenarios, such as question answering (Gu et al., 2023), information retrieval (Ren et al., 2023) and recommendation systems (Liu et al., 2023). However, LLMs have tendency to generate non-existing facts when faced with questions that are out of their parametric knowledge (Maynez et al., 2020). Many efforts are dedicated to mitigating hallucinations in LLMs, such as retrieval-augmented generation (Gao et al., 2024,Peng et al., 2023), multi-agent debate (Du et al., 2023,Sun et al., 2023), and model confidence calibration (Zhang et al., 2024, Hu et al., 2023).
Knowledge Boundary: There are many different ways to utilize knowledge boundary to reduce LLM hallucination. Liang et al. (2024b)’s work uses merged knowledge probing and consistency checking methods to help LLM express their internal knowledge. Chen et al. (2024)’s work leverages LLM internal signals to let LLM know their unknowns. Zhang et al. (2024) utilize knowledge boundary to instruct LLM say "I don’t know". It is a popular way to use confidence to express knowledge boundary of LLMs and we also follow this.
Multiple Problem Setting: Current LLM research has predominantly focused on single problem setting. There are only a few works focusing on this new setting. (Cheng et al., 2023) propose batch prompting that prompts LLMs with single independent problems batched together following few-shot exemplars together. Son et al. (2024) goes further by researching sequential datasets and develops the first multi-task benchmark (MTI-Bench). Wang et al. (2024) pays attention to zero-shot cases of multi-problem setting and design a new benchmark ZeMPEB. Overall, LLM hallucination under this multi-problem setting is still under-explored.
A.3 Template for QA-Confidence pair
Question: <Question>. Answer: <Answer>. Are you sure you accurately answered the question based on your internal knowledge? 1: <Confidence> 2: <Confidence> 3: <Confidence>
A.4 Dataset Details
We carry out our experiments across six datasets, described as follows.
  • GSM (Cobbe et al., 2021): a dataset containing high-quality grade school math problems created by the OpenAI group. These problems require between 2 and 8 steps to solve, primarily involving a sequence of elementary calculations with basic arithmetic operations such as addition, subtraction, multiplication, and division to arrive at the final answer. We directly use 7.5k training data and 1k testing data in our Question Answer setting.
  • Pararel (Elazar et al., 2021): a dataset containing factual knowledge with a variety of prompts and relationships, originally created for mask prediction. In Question Answer setting, we employ the modified dataset from Zhang et al. (2024).
  • MMLU (Hendrycks et al., 2021) a dataset covering different subjects and difficulty. It tests both world knowledge and problem solving ability, which has good granularity and breath. We directly use the modified dataset from Zhang et al. (2024) in our Multiple Choice setting.
  • CoQA (Reddy et al., 2019) a dataset designed to evaluate the ability of models to understand and generate answers in a conversational setting. We randomly pick 5k training dataset from theirs. In Question Answer setting, we combine multiple questions together under the same "story" category in the dataset.
  • MTI Bench (Son et al., 2024) a comprehensive evaluation benchmark encompassing 5,000 instances across 25 tasks. We pick the sequential part of this benchmark and divide it into 800 training data and 200 test data.
  • SQA (Iyyer et al., 2017) a dataset designed to explore the task of answering sequences of inter-related questions on HTML tables. We pick 5 sequential questions for each HTML table and have 3985 training data.
A.5 Formula and Calculation Details
Average Precision (AP) Score measures the performance of a binary classifier’s confidence rankings. It corresponds to the area under the Precision-Recall curve. It is calculated as follows:
$\mathrm{AP}=\sum_{k=1}^{n}\Bigl{(}R_{k}-R_{k-1}\Bigr{)}\times P_{k}$
where k is the number of data at current thread with precision $P_{k}$ and recall $R_{k}$. n is the total data number. The confidence is the weighted average of certain prediction probability and uncertain prediction probability.
Expected Calibrated Error (ECE) indicates how well a model’s predicted probabilities match the true likelihood of an event. We split the predictions into 10 bins based on the certain prediction probability, then compare the average predicted probability to the actual proportion of positive samples (correct cases) in each bin. It is calculated as follows:
$\mathrm{ECE}=\sum_{m=1}^{10}\frac{\lvert B_{m}\rvert}{n}\bigl{\lvert}\overline{p}_{m}-\overline{y}_{m}\bigr{\rvert}$
where m is the bin number with corresponding average predicted probability $\overline{p}_{m}$ and actual proportion of positive samples $\overline{y}_{m}$.
A.6 Implementation
We use HuggingFace PEFT (Mangrulkar et al., 2022) to conduct LoRA fine-tuning (Hu et al., 2021). We set the training epoch to 3, learning rate to $1\mathbf{e}^{-5}$, LoRa rank to 8, and LoRa scaling factor to 32. The batch size is 1 and the temperature is 0. All experiments are implemented on Nvidia A100-40GB GPUs.
A.7 Case Study
We show two specific cases for MAC-Tuning under the multiple problem setting with question number n = 3 in Figure 5. The example on the left is from the SQA dataset, in which a table context is given and the LLM need to answer sequential questions based on the table. LLM answers correctly and shows certainty to first two questions, so these two questions will be counted into accuracy calculation. It answers wrong and shows uncertainty to the third question, which achieves the refusal behavior that we aim to see. The example on the right is from the GSM dataset. The LLM gives wrong answers to the second question but indicates certainty, which means this is a failure case.
Figure 5: Specific cases for MAC-Tuning under the multiple problem setting with question number n = 3. Red-highlighted context indicates inaccuracies in the LLM’s output. The left example is drawn from the Sequential setting dataset (SQA), while the right example is from the Independent setting dataset (GSM), with one-shot context omitted for conciseness.
A.8 Detailed Information for Variant Methods
The detailed example for different baseline methods is shown in Figure 6.
Figure 6: A specific case to show how baseline methods are doing the fine-tuning. The answers are derived from the highlighted portions of the context. In QA-Only, the input is the Question instruction, and the output is the Answer. In Merge-AC, the output includes both the Answer and its Confidence. Single-QA is the single-problem variant of Merge-AC.
A.9 Cross Task Transfer Study
We fine-tune the model with question number n = 3 and let it response to question number n = 1 (which is Single Problem setting) and question number n = 5. For the former one, we want to test if the model understands the single problem. For the latter one, we want to test if MAC-Tuning can generalize across different question number n. The results are reported in Table 5.
Question Number CoQA ParaRel GSM MMLU
n = 1 78.78 84.24 71.12 54.60
n = 5 79.12 86.16 67.67 63.69
Table 5: Accuracy (%) for MAC-Tuning with question number n = 3 transferring to question number n = 1 (which is single-problem setting) and question number n = 5. We use one-shot CoT for GSM results.
From the result of n = 1, we observe that accuracy improves in easy dataset like CoQA but decreases in difficult dataset like GSM, comparing with using question number n = 1 to fine-tune. This indicates that LLM learn knowledge during fine-tuning rather than just memorizing the patterns for multi-problem setting. With n = 5, the model performs about the same or even better than using question number n = 5 to fine-tune. This strengthen the statement we make in Section 3.6 that LLMs are more efficient at learning multiple easy tasks but struggle with several difficult tasks simultaneously.
A.10 Certainty Distribution of the Training Dataset
We demonstrate the certainty distribution of the training dataset under Multiple Problem setting with question number n = 3 in Figure 7:
Figure 7: Certainty distribution of the training set under multi-problem setting with n = 3